# XAL: EXplainable Active Learning Makes Classifiers Better Low-resource Learners

### 14 Mar 2024 v2

@@ 2 Related Work
Large Language Model. Recently, LLMs of generative schema have shown excellent performance in various NLP tasks. However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021). Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks. Some studies distill the knowledge of large language models to a smaller one Hsieh et al. (2023) and some also use LLMs to augment data to achieve stronger text classification performance Ye et al. (2022); Yu et al. (2023). LLMs has also demonstrated a strong capability in generating high-quality reasoning steps Hsieh et al. (2023); Wei et al. (2022); Kojima et al. (2022). In this study we aim to distill the reasoning ability of LLMs to smaller models, and encourage the models to distinguish the reasonablity of explanations to identify informative data in AL scenario.
@@ 3.4 Generation of Golden Explanations
Previous work has shown that LLMs are good at reasoning (Bang et al., 2023; Rajasekharan et al., 2023). Inspired by these studies, we take the LLMs, such as ChatGPT and GPT4, as the teacher models, and query them to generate explanations for each selected labeled data, eliminating the annotation cost of human labor. In particular, we design slightly different prompt templates for different tasks, and the prompt for each task is shown in Appendix A. Taking stance detection as an example, its prompt template is designed as ‘The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words’, where the Target is the corresponding stance target, and the Label is the classification label. The final query to the teacher model is the concatenation of the text and the prompt. We construct a reasonable explanation by feeding the golden label into the query and generate several unreasonable explanations by feeding wrong labels. Figure 7 shows an example that we generate explanations by querying ChatGPT, where we can observe that ChatGPT could provide different explanations according to the given label.
@@ 4.1 Tasks and Dataset
We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text; (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff; (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target; (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment; (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic. We adopt the datasets RTE Bentivogli et al. (2009), MRPC Dolan et al. (2004), MAMS Jiang et al. (2019), COVID19 Glandt et al. (2021), DEBA Pougué-Biyong et al. (2021) and CLEF Kanoulas et al. (2017) for each task, respectively. Appendix A demonstrates the details of six datasets with examples.  [^1] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1. In DEBA, we adopt the topic of climate change for experiments.
@@ Appendix A Tasks and Corresponding Prompts
We first introduce the dataset and its statistics in Table 3. We show the tasks and examples for experiments in Table 4, including natural language inference, paraphrase detection, category sentiment classification, stance detection, (dis)agreement detection, and relevance classification. Then we show how we obtain different explanations in Figure 7 with an example of COVID19. We also show the prompts we used for explanation generation through querying ChatGPT (Table 5).
@@
Task
Dataset
# Labels Train Dev Test
Natural Language Inference
RTE (Bentivogli et al., 2009)
2 2,240 250 278
Paraphrase Detection
MRPC (Dolan et al., 2004)
2 3,667 409 1,726
Stance Detection
COVID19 (Glandt et al., 2021)
3 4,533 800 800
Category Sentiment Classification
MAMS (Jiang et al., 2019)
3 7,090 888 901
(Dis)agreement Detection
DEBA (Pougué-Biyong et al., 2021)
3 4,617 578 580
Relevance Classification
CLEF (Kanoulas et al., 2017)
2 7,847 981 982

### 15 Mar 2024 v3

@@ 2 Related Work
Large Language Model. Recently, LLMs of generative schema have shown excellent performance in various NLP tasks. However, some studies show that in-context learning based on LLMs (Radford et al., 2019; Brown et al., 2020) suffers from practical issues such as high computation costs for inference (Liu et al., 2022), inclination to their internal knowledge Yan et al. (2024), catastrophic forgetting during instruction tuning Luo et al. (2023b), over-sensitive to example choices and instruction wording (Gao et al., 2020; Schick and Schütze, 2021). Considering the problems, LLMs are also applied to improve the generalization of smaller models for specific tasks. Some studies distill the knowledge of large language models to a smaller one Hsieh et al. (2023) and some also use LLMs to augment data to achieve stronger text classification performance Ye et al. (2022); Yu et al. (2023). LLMs has also demonstrated a strong capability in generating high-quality reasoning steps Hsieh et al. (2023); Wei et al. (2022); Kojima et al. (2022). In this study we aim to distill the reasoning ability of LLMs to smaller models, and encourage the models to distinguish the reasonablity of explanations to identify informative data in AL scenario.
@@ 3.4 Generation of Golden Explanations
Previous work has shown that LLMs are good at reasoning (Bang et al., 2023; Rajasekharan et al., 2023; Hsieh et al., 2023). Inspired by these studies, we take the LLMs, such as ChatGPT and GPT4, as the teacher models, and query them to generate explanations for each selected labeled data, eliminating the annotation cost of human labor. In particular, we design slightly different prompt templates for different tasks, and the prompt for each task is shown in Appendix A. Taking stance detection as an example, its prompt template is designed as ‘The stance of this tweet to the target {Target} is {Label}, explain the reason within 50 words’, where the Target is the corresponding stance target, and the Label is the classification label. The final query to the teacher model is the concatenation of the text and the prompt. We construct a reasonable explanation by feeding the golden label into the query and generate several unreasonable explanations by feeding wrong labels. Figure 7 shows an example that we generate explanations by querying ChatGPT, where we can observe that ChatGPT could provide different explanations according to the given label.
@@ 4 Experiments
Task
Dataset
# Labels Train Dev Test
Natural Language Inference
RTE (Bentivogli et al., 2009)
2 2,240 250 278
Paraphrase Detection
MRPC (Dolan et al., 2004)
2 3,667 409 1,726
Stance Detection
COVID19 (Glandt et al., 2021)
3 4,533 800 800
Category Sentiment Classification
MAMS (Jiang et al., 2019)
3 7,090 888 901
(Dis)agreement Detection
DEBA (Pougué-Biyong et al., 2021)
3 4,617 578 580
Relevance Classification
CLEF (Kanoulas et al., 2017)
2 7,847 981 982
@@ 4.1 Tasks and Dataset
We conduct experiments on six different text classification tasks: (1) Natural Language Inference aims to detect whether the meaning of one text is entailed (can be inferred) from the other text Bentivogli et al. (2009); Wang et al. (2018); (2) Paraphrase Detection requires identifying whether each sequence pair is paraphrased; (3) Category Sentiment Classification aims to identify the sentiment (Positive/Negative/Neutral) of a given review to a category of the target such as food and staff Jiang et al. (2019); Luo et al. (2022a); (4) Stance Detection aims to identify the stance (Favor/Against/Neutral) of a given text to a target Glandt et al. (2021); Luo et al. (2022b); (5) (Dis)agreement Detection aims to detect the stance (Agree/Disagree/Neutral) of one reply to a comment Pougué-Biyong et al. (2021); Luo et al. (2023a); (6) Relevance Classification aims to detect whether a scientific document is relevant to a given topic Kanoulas et al. (2017); Guo et al. (2023). Appendix A demonstrates the details of six datasets with examples.  [^1] Without losing generality, we randomly split the training set in RTE, and MRPC into train/dev set with proportion 9:1. In DEBA, we adopt the topic of climate change for experiments.
@@ Appendix A Tasks and Corresponding Prompts
We show the tasks and examples for experiments in Table 4, including natural language inference, paraphrase detection, category sentiment classification, stance detection, (dis)agreement detection, and relevance classification. Then we show how we obtain different explanations in Figure 7 with an example of COVID19. We also show the prompts we used for explanation generation through querying ChatGPT (Table 5).